# The Error in Rayleigh’s Approximative Period

### v3

@@ 1 Introduction
Therefore we offer the following theorem to fill this gap.
@@ 3 Upper and Lower Bounds for the True Period
The exact formula for the period gives $34.1941$ seconds while Rayleigh’s approximative period gives $35.4807$ seconds. Therefore the absolute value of the relative error is $3.4\%$. The bounds in (1.8), after we subtract $1$ give 2.5%⩽3.4%⩽5% which verifies the bounds in the theorem in this case.
@@
Our bounds on the relative error permit us to explain why Rayleigh’s approximation is so bad in this case.
If we use the Taylor expansion of each bound we find
$\frac{1}{2\left(\frac{L}{L_{0}}-1\right)}\left(\frac{y_{0}}{2L}\right)^{2}+O\left(\frac{y_{0}}{2L}\right)^{4}\leqslant|R|\leqslant\frac{1}{\left(\frac{L}{L_{0}}-1\right)}\left(\frac{y_{0}}{2L}\right)^{2}+O\left(\frac{y_{0}}{2L}\right)^{4}$
which shows that to a first approximation, the relative error is directly proportional to the square of the initial fractional displacement and inversely proportional to the initial stretch. Thus, for a given displacement, $\frac{y_{0}}{2L}$, if the stretch is tiny, that is if $L\approx L_{0}$, the relative error will be very large. In this particular case the stretch is $\frac{1}{1570}=0.00064$ meters, and thus the lower bound’s leading term is $15.7\%$ which shows that the true relative error will be even larger. And, indeed, above we found it to be about $25\%$.

### v4

@@ 1 Introduction
Therefore we offer the following theorems to fill this gap.
@@
We can combine these radical function bounds into a simple and elegant formula for the (absolute value of) the relative error which displays is true dependence on the various variables in play.
Theorem 2.
$\boxed{|R|=\Theta\cdot\frac{L_{0}}{L-L_{0}}\cdot\left(\frac{y_{0}}{2L}\right)^{2}}$ (1.9)
where
$\frac{\frac{1}{2}-\left(\frac{y_{0}}{2L}\right)^{2}}{1+\frac{1}{32}\frac{L_{0}}{L-L_{0}}}\leqslant\Theta\leqslant 1.$ (1.10)
So, the (absolute value of) the relative error is directly proportional to the square of the initial fractional displacement $\left(\frac{y_{0}}{2L}\right)^{2}$ and inversely proportional to the initial stretch $L-L_{0}$. The factor of proportionality $\Theta$ is always greater than $\frac{\frac{1}{2}-\left(\frac{y_{0}}{2L}\right)^{2}}{1+\frac{1}{32}\frac{L_{0}}{L-L_{0}}}$ but always is less than $1$ .
Thus, notwithstanding the analytical complexity of the radical function bounds, their behavior is limited by the simple bounds (1.10) on $\Theta$.
@@ 4 Upper and lower bounds on the relative error
5 Proof of the formula for the relative error.
First we prove the upper bound.
$|R|\leqslant\left(\frac{y_{0}}{2L}\right)^{2}$ (5.1)
By the right-hand side of (3.2)
$|R|$ $\leqslant$ $\sqrt{\frac{\frac{1}{L_{0}}-\frac{1}{\sqrt{L^{2}+y_{0}^{2}}}}{\frac{1}{L_{0}}-\frac{1}{L}}}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{1}{L_{0}}-\frac{1}{\sqrt{L^{2}+y_{0}^{2}}}}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{1}{L_{0}}-\frac{1}{L}+\frac{1}{L}-\frac{1}{\sqrt{L^{2}+y_{0}^{2}}}}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{1}{L_{0}}-\frac{1}{L}+\frac{1}{L}-\frac{1}{L}+\frac{y_{0}^{2}}{2L^{3}}-\frac{3y_{0}^{4}}{8L^{5}}+-\cdots}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{L-L_{0}}{L-L_{0}}+\frac{y_{0}^{2}}{2L^{3}}-\frac{3y_{0}^{4}}{8L^{5}}+-\cdots}-1$
$=$ $\sqrt{1+\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{2L^{2}}-\frac{3y_{0}^{4}}{8L^{4}}+-\cdots\right)}-1$
$<$ $1+\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{3y_{0}^{4}}{16L^{4}}+-\cdots\right)-1$
$<$ $\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}\right)$
where we have used the fact that a convergent alternating series whose terms tend monotonically to zero has a sum that is between two consecutive partial sums and that $\sqrt{1+u}<1+\frac{u}{2}.$ This completes the proof of the upper bound.
Now we prove the lower bound.
$|R|\geqslant\frac{\frac{1}{2}-\left(\frac{y_{0}}{2L}\right)^{2}}{1+\frac{1}{32}\frac{L_{0}}{L-L_{0}}}\cdot\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}}{2L}\right)^{2}.$ (5.2)
By the right-hand side of (3.2)
$|R|$ $\geqslant$ $\sqrt{\frac{\frac{1}{L_{0}}-\frac{2}{L+\sqrt{L^{2}+y_{0}^{2}}}}{\frac{1}{L_{0}}-\frac{1}{L}}}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{1}{L_{0}}-\frac{2}{L+\sqrt{L^{2}+y_{0}^{2}}}}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{1}{L_{0}}-\frac{1}{L}+\frac{1}{L}-\frac{2}{L+\sqrt{L^{2}+y_{0}^{2}}}}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{1}{L_{0}}-\frac{1}{L}+\frac{1}{L}-\frac{1}{L}+\frac{y_{0}^{2}}{4L^{3}}-\frac{y_{0}^{4}}{8L^{5}}+-\cdots}-1$
$=$ $\sqrt{\frac{LL_{0}}{L-L_{0}}}\cdot\sqrt{\frac{L-L_{0}}{L-L_{0}}+\frac{y_{0}^{2}}{4L^{3}}-\frac{y_{0}^{4}}{8L^{5}}+-\cdots}-1$
$=$ $\sqrt{1+\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}+-\cdots\right)}-1$
$>$ $\sqrt{1+\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}\right)}-1$
$>$ $1+\frac{\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}\right)}{2+\frac{\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}\right)}{2}}-1$
$=$ $\frac{\frac{1}{2}\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}\right)}{1+\frac{\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}\right)}{4}}$
$=$ $\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}}{2L}\right)^{2}\frac{\left(\frac{1}{2}-\left(\frac{y_{0}}{2L}\right)^{2}\right)}{1+\frac{\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}\right)}{4}}$
where we used the inequality $\sqrt{1+u}>1+\frac{u}{2+\frac{u}{2}}.$. Now the function $\frac{y_{0}^{2}}{4L^{2}}-\frac{y_{0}^{4}}{8L^{4}}$ is increasing for $0\leqslant y_{0}\leqslant L$ which means that it takes it maximum value $\frac{1}{8}$ when $y_{0}=L.$ Plugging this into the denominator of the last fraction we obtain
$|R|>\frac{L_{0}}{L-L_{0}}\left(\frac{y_{0}}{2L}\right)^{2}\frac{\left(\frac{1}{2}-\left(\frac{y_{0}}{2L}\right)^{2}\right)}{1+\frac{1}{32}\frac{L_{0}}{L-L_{0}}}\equiv\frac{L_{0}}{L-\frac{31}{32}L_{0}}\left(\frac{y_{0}}{2L}\right)^{2}\left(\frac{1}{2}-\left(\frac{y_{0}}{2L}\right)^{2}\right)$ (5.3)
The second formula in (5.3) is an equally elegant form as the original.
This completes the proof of the formula for the relative error.
6 Examples
@@
The exact formula for the period gives $34.1941$ seconds while Rayleigh’s approximative period gives $35.4807$ seconds. Therefore the absolute value of the relative error is $3.4\%$. The bounds in (1.10), give 2.5%⩽3.4%⩽5% which verifies the bounds in the theorem in this case.
@@
Our bounds on the relative error permit us to explain why Rayleigh’s approximation is so bad in this case. Namely, for a given displacement, $\frac{y_{0}}{2L}$, if the stretch is tiny, that is if $L\approx L_{0}$, the relative error will be very large. In this particular case the stretch is $\frac{1}{1570}=0.00064$ meters, and thus the lower bound is $15.7\%$ which shows that the true relative error will be even larger. And, indeed, above we found it to be about $25\%$.